OpenStudy (anonymous):

lim x -> infinity [sqrt(4x^2 - 3x+ 5) - 2x]. Please help me with the steps I would need to go through to solve this.

OpenStudy (anonymous):

Is it zero?

OpenStudy (anonymous):

what do u think about my solution?

OpenStudy (anonymous):

why did you just divide by x^2?

OpenStudy (anonymous):

We need help @Hero . I am out of the words. My English is weak

OpenStudy (anonymous):

yes my solution is incorrect sorry

OpenStudy (zarkon):

multiply by \[\frac{\sqrt{4x^2 - 3x+ 5} + 2x}{\sqrt{4x^2 - 3x+ 5} + 2x}\]

OpenStudy (anonymous):

thanks. but then i still have \[(-3x+5)/(\sqrt(4x^2 - 3x + 5) + 2x)\]

OpenStudy (anonymous):

what do i do from there?

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

dear when u get the answer u wrote above u can put infinitive instead x and u can ignore all other term except -3x /squr (4x^2) +2x so u will have -3x / 4x and so u have -3/4 .

OpenStudy (anonymous):

I dont understand what you just said. sorry.

OpenStudy (anonymous):

let me draw what i mean !

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sri just read the green part .

OpenStudy (anonymous):

when you have infinitive because it's so great u can just use the term with the greatest power in num and dem . and u can ignore the other part or u can eliminate them !

OpenStudy (anonymous):

did u get it ? i u didn't get it i can explain it more

OpenStudy (anonymous):

thanks. i think i understand now :)

OpenStudy (anonymous):

ok dear , hope u be successful in ur exams .

OpenStudy (zarkon):

I would factor out an x from the top and bottom